### v1

# Atomic Calibration of LLMs in Long-Form Generations

Abstract
Large language models (LLMs) often suffer from hallucinations, posing significant challenges for real-world applications. Confidence calibration, which estimates the underlying uncertainty of model predictions, is essential to enhance the LLMs’ trustworthiness. Existing research on LLM calibration has primarily focused on short-form tasks, providing a single confidence score at the response level (macro calibration). However, this approach is insufficient for long-form generations, where responses often contain more complex statements and may include both accurate and inaccurate information. Therefore, we introduce atomic calibration, a novel approach that evaluates factuality calibration at a fine-grained level by breaking down long responses into atomic claims. We classify confidence elicitation methods into discriminative and generative types and demonstrate that their combination can enhance calibration. Our extensive experiments on various LLMs and datasets show that atomic calibration is well-suited for long-form generation and can also improve macro calibration results. Additionally, atomic calibration reveals insightful patterns in LLM confidence throughout the generation process.
Atomic Calibration of LLMs in Long-Form Generations
Caiqi Zhang♥†, Ruihan Yang♠†, Zhisong Zhangthanks: Corresponding authors., Xinting Huang, Sen Yang♣†, Dong Yu, Nigel Collier11footnotemark: 1 University of Cambridge, Fudan University, Tencent AI Lab, The Chinese University of Hong Kong {cz391, nhc30}@cam.ac.uk, zhisonzhang@tencent.com
footnotetext: Work done during the internship at Tencent AI Lab.
1 Introduction
Figure 1: Comparison between traditional macro calibration in response-level and our atomic calibration. The Fact. label is assigned by fact-checking module. We only list three atomic claims for illustration.
While large language models (LLMs), such as Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023), and GPT models (OpenAI, 2022), excel in various tasks, they still struggle with faithfulness and reliability issues. LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023). To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017). A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024).
Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024). These studies often use datasets like TriviaQA (Joshi et al., 2017) and Natural Questions (Joshi et al., 2017), where answers typically have fewer than 10 words. However, in real-world applications, responses to user queries are often much longer (Zhang et al., 2024), sometimes extending to hundreds or even thousands of words. In such cases, the quality of LLM responses is not simply binary (correct or incorrect), as answers may include both accurate and inaccurate statements. Previous work on long-form calibration has primarily focused on response-level calibration (Zhang et al., 2024; Huang et al., 2024) (which we term macro calibration). This approach provides an overall confidence estimation for the entire response (as shown in the upper part of Figure 1). However, it is insufficient for long-form generation, as it cannot adequately capture the model’s fine-grained uncertainty across multiple factual statements.
In this work, we introduce the concept of atomic calibration, which evaluates calibration at the fine-grained level of atomic claims (as illustrated in the lower part of Figure 1). We focus on the aspect of factuality, since hallucination are a widely recognized issue of LLMs (Zhang et al., 2023b; Huang et al., 2023) and the factuality of statements can be objectively determined. Accurate factual calibration is crucial for mitigating the hallucination problem (Mahaut et al., 2024). To perform atomic calibration, we decompose long responses into self-contained atomic claims and evaluate the calibration for all atomic claims. Various confidence elicitation methods are studied to provide accurate confidence estimation. We provide a detailed investigation into long-form generation calibration, covering 7 LLMs and 3 datasets. The main contributions of our work are as follows:
  • We propose the novel concept of atomic calibration. Unlike macro calibration, which provides a single response-level confidence score, atomic calibration offers a more fine-grained analysis of model calibration, making it better suited for long-form generation (§2).
  • We introduce a categorization of confidence elicitation methods for long-form generation into two types: discriminative and generative, estimating the model’s intrinsic and external confidence, respectively. We examine five confidence elicitation methods and assess their effectiveness for atomic calibration (§3).
  • We show that atomic calibration is particularly well-suited for long-form generation, and that macro calibration can be enhanced by incorporating atomic calibration results. We also find the two types of confidence elicitation methods are complementary and their fusion can provide better calibration results. Atomic calibration also enables deeper analysis, revealing patterns in confidence and calibration changes throughout the generation process (§5).
2 Atomic Calibration
For a language model $\mathcal{M}$, let $x\in\mathcal{X}$ represent the response generated by model $\mathcal{M}$ for a query $q$, and let $y\in\mathcal{Y}_{t}$ denote the corresponding label, where $\mathcal{Y}_{t}\subseteq[0,1]$ indicates a quality score for a specific task $t\in T$. Note that unlike multiple-choice and short-form questions, which typically only focus on the correctness of the answer, the tasks in $T$ involve various aspects, such as factuality, coherence, creativity, etc.
We define a probability prediction function $f:\mathcal{X}\to\Delta^{\lvert\mathcal{Y}_{t}\rvert}$, where $\Delta^{\lvert\mathcal{Y}_{t}\rvert}$ represents the $\lvert\mathcal{Y}_{t}\rvert$-dimensional simplex. In this context, $f(x)_{y}$ denotes the probability assigned to the label $y$ given the generated output $x$. In the remaining of this work, we focus specifically on calibrating for factuality. In this case, $\mathcal{Y}$ represent $\mathcal{Y}_{t}$ when $t$ corresponds to factuality. $\mathcal{Y}$ takes values in [0, 1], indicating the factuality level of a response. Based on the calibration concepts discussed in (Guo et al., 2017), we define the calibration of each response as follows:
Definition 1 (Macro Calibration on Factuality)
A language model $\mathcal{M}$ that produces generations $x\sim\mathcal{M}(x\mid q)$ is said to be response-level (macro) calibrated if
$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta,\quad\forall\beta\in\Delta^{\lvert\mathcal{Y}\rvert}.$
In the context of long-form generation, a single response $x$ may encompass multiple atomic claims. Macro calibration at the response level cannot fully present the fine-grained uncertainty at the atomic level. To address this, we decompose the response $x$ into $N$ atomic claims $c_{i}$, represented as $x=\coprod_{i=1}^{N}c_{i}$. Each atomic claim $c_{i}$ is assigned a binary label $y_{i}\in\mathcal{Y}_{i}$, where $\mathcal{Y}_{i}=\{0,1\}$, indicating its truthfulness. The overall factuality score for the response $y$ is computed as $y=\frac{1}{N}\sum_{i=1}^{N}y_{i}$. Similarly, we define $f(c_{i})_{y_{i}}$ as the probability of the label $y_{i}$ given the atomic claim $c_{i}$. Building on this decomposition, we propose a fine-grained measure of calibration at the atomic level as follows:
Definition 2 (Atomic Calibration on Factuality)
A language model $\mathcal{M}$, which generates a long-form response $x$ conditioned on the query $q$, $x\sim\mathcal{M}(x\mid q)$, is considered atomic-level calibrated if, for each atomic claim $c_{i}$ with its corresponding label $y_{i}$, the following condition holds:
$\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta_{i},\quad\forall\beta_{i}\in\Delta^{\lvert\mathcal{Y}_{i}\rvert}.$
Remarks:
(1) Unlike traditional classification problems where $f(x)_{y}$ is usually represented as a single log probability of the predicted answer, it is much more challenging to measure model confidence in text generation tasks. Different confidence elicitation methods may yield different predictions of the $f(x)_{y}$; therefore, how to design proper elicitation methods is a key problem. (2) Macro calibration is not equivalent to the sum of atomic calibrations, as illustrated by:
$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta$
⇏$\not\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta$
⇏$\not\Rightarrow$ $\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta,\forall i\in\{1,....,N\}.$
3 Confidence Elicitation Methods
To calculate the confidence scores for model outputs in long-form generation, we define two types of confidence elicitation methods: generative and discriminative. Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty. The more frequently certain facts are covered by the samples, the higher the model’s confidence in those facts. In contrast, discriminative methods assess uncertainties by asking the model itself. This is motivated by the findings that models tend to perform better on discriminative tasks (Saunders et al., 2022), and thus they may already possess the capability to estimate the confidence of their own outputs in a discriminative manner.
We first generate one response $x$ as the answer to the query $q$. Then, $x$ is broken into atomic claims $C$. Following (Min et al., 2023; Wei et al., 2024; Zhao et al., 2024), each atomic claim contains a single piece of information and must be self-contained. For generative methods, we sample an additional set of responses $K$, and compare them against the original response $x$. For each atomic claim in $C$, we assign it a confidence score.
3.1 Generative Methods
Gen-Binary. The basic assumption here is that if a fact is frequently conveyed when sampled multiple times, the model is considered “confident” about that fact. For an atomic claim $c_{i}$ in $C$, we utilize a natural language inference model $\mathcal{M}_{\mathrm{NLI}}$ to examine whether $c_{i}$ is supported or not supported by each of the additional samples. Let $K_{s}$ be the set of samples supporting $c_{i}$. Then, the confidence in $c_{i}$ is calculated as:
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K\rvert}$
Gen-Multi. Gen-Multi assumes that the model is more confident in facts that are consistently expressed. Unlike Gen-Binary, it further divides the “not supported” class $K_{ns}$ into “conflict” ($K_{c}$) if the fact is presented differently in the sample, and “not mentioned” ($K_{nm}$) if the fact is not mentioned in the sample. We then calculate the confidence by only considering supporting and conflicting samples:
$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K_{s}\rvert+\lvert K_{c}\rvert}$
3.2 Discriminative Methods
Dis-Single. Following (Kadavath et al., 2022; Tian et al., 2023), we directly ask the model whether one single atomic claim is true or false. The probability the model assigns to token “True” ($P(true)$) in its generation is viewed as the confidence. As each atomic claim is judged individually, one advantage of this method is that there is no cross-claim influences when the model makes confidence judgments.
Dis-Context. In addition to the method where each claim is judged in a self-contained way, we also consider a setting where additional context is provided. Here, the context denotes the passage where the atomic claim is extracted, or the prompt that generates the response. The context helps the model to more accurately locate the atomic claim, and thus potentially leads to better confidence elicitation. $P(true)$, given the context, is then used as the confidence score, just as in Dis-Single.
Dis-Rating. Instead of using $P(true)$, in Dis-Rating, we directly prompt the model to assign a numerical value representing its confidence in the atomic claim $c_{i}$. A score of 0 indicates no confidence, while 10 represents maximum confidence. An alternative approach is to use semantic expressions ranging from “Very Uncertain” to “Very Confident”. However, Tian et al. (2023) show that LLMs express uncertainty as effectively, or even more effectively, using numerical values rather than words.
3.3 Confidence Fusion Strategies
Since the generative and discriminative methods provide different perspectives of confidence estimation, it is natural to explore fusion methods that combine different confidence estimates for better calibration. After exploration, we consider four simple but effective options: (1) MinConf selects the minimum of the two confidence values, providing a conservative estimate; (2) HMean calculates the harmonic mean, providing a balanced fusion that weights lower confidence values more heavily, to help prevent overconfident predictions; (3) ProdConf multiplies the confidences, ensuring a high final confidence only when both estimates are strong, thereby effectively penalizing discrepancies between the two estimates; and (4) WAvg computes the weighted average, with a hyper-parameter weight, prioritizing the more reliable estimate and allowing flexibility to adjust the influence of each confidence source based on prior knowledge or validation results.
4 Evaluation Metrics
4.1 Atomic Calibration
In atomic calibration, we utilize FActScore (Min et al., 2023) to measure the atomic claims’ factuality. For each claim, we have its binary factuality label and a continuous confidence score. Since this corresponds to standard calibration scenarios, we follow the conventions from previous work (Kuhn et al., 2022; Zhu et al., 2023; Tian et al., 2023) and use Expected Calibration Error (ECE) (Naeini et al., 2015), Brier Score (Brier, 1950), and AUROC to assess calibration. Details of the three metrics are in Appendix A.
4.2 Macro Calibration
In macro calibration, factuality scores for responses are represented as continuous values from 0 to 1, therefore, the atomic calibration metrics mentioned above are not directly applicable. Apart from the Spearman Correlation, we propose two novel metrics UCCE and QCCE that extend ECE to continuous factuality scores.
Spearman Correlation. Following (Zhang et al., 2024; Huang et al., 2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores. Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
Uniform Continuous Calibration Error (UCCE). UCCE evaluates how well a model’s continuous predictions align with the true values by dividing the predicted values into bins of equal size. This metric assesses the model’s calibration across the range of predictions. The calibration error is computed by comparing the mean predicted and true values within each bin.
Let $\hat{y}_{i}$ represent the predicted value for the $i$-th sample, $y_{i}$ the true value for the $i$-th sample, $B_{m}$ the set of samples falling into the $m$-th bin, $M$ the number of bins, and $N$ the total number of samples. The UCCE is computed as follows:
$\text{UCCE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\frac{\sum_{i\in B_{m}}\hat{y}_{i}}{|B_{m}|}-\frac{\sum_{i\in B_{m}}y_{i}}{|B_{m}|}\right|.$
UCCE is particularly effective for scenarios where predicted values are uniformly distributed.
Quantile Continuous Calibration Error (QCCE). QCCE uses a similar approach to UCCE but utilizes quantile binning, ensuring that each bin contains an equal number of samples. Consequently, each bin has a size of $\lvert B_{m}\rvert=\frac{N}{M}$. QCCE is advantageous for skewed or non-uniform distributions of predicted values, as it avoids sparsity in any bin by ensuring a balanced representation across all bins. The QCCE is calculated as follows:
QCCE $=\frac{1}{M}\sum_{m=1}^{M}\left|\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{y}_{i}-\frac{1}{|B_{m}|}\sum_{i\in B_{m}}y_{i}\right|$
$=\frac{1}{N}\sum_{m=1}^{M}\left|\sum_{i\in B_{m}}\hat{y}_{i}-\sum_{i\in B_{m}}y_{i}\right|.$
Bios LongFact WildHallu
ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
Llama3-8B-Instruct
Dis-Context 35.5 35.8 74.5 11.9 13.6 74.4 12.5 16.5 83.5
Dis-Rating 26.8 29.0 71.1 3.5 12.0 66.9 5.3 15.2 79.8
Dis-Single 32.6 33.9 74.5 14.3 15.2 69.8 19.2 20.9 79.3
Gen-Binary 10.0 17.8 83.1 8.5 11.4 77.3 11.1 15.2 82.0
Gen-Multi 37.4 37.3 64.2 12.6 13.1 58.5 21.9 22.1 65.4
Mistral-7B-Instruct
Dis-Context 24.8 26.0 77.5 15.7 16.1 75.3 20.6 21.7 79.8
Dis-Rating 44.5 42.5 65.0 10.0 14.2 67.9 19.7 23.9 68.1
Dis-Single 30.2 30.7 75.2 20.4 20.5 66.6 24.0 24.6 75.1
Gen-Binary 13.7 19.0 81.9 8.4 11.5 80.1 12.7 17.0 81.3
Gen-Multi 42.2 41.8 65.0 13.4 13.9 61.7 26.6 26.4 64.2
Qwen2-7B-Instruct
Dis-Context 26.5 28.3 75.5 13.9 14.8 77.9 17.2 19.4 81.2
Dis-Rating 41.5 39.7 64.2 3.5 11.7 62.6 8.2 18.1 70.4
Dis-Single 29.3 30.4 75.5 16.1 16.8 74.7 18.7 20.3 80.1
Gen-Binary 10.9 16.7 83.8 6.3 9.9 81.9 9.5 14.0 82.5
Gen-Multi 41.7 41.1 65.6 11.6 12.1 62.8 21.0 21.0 64.4
Table 1: Atomic Calibration Results. All the numbers are in percentages.
5 Experiments and Results
5.1 Experiment Setup
Models. We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024).
Datasets. We use three datasets for long-form QA: Bios (Min et al., 2023), which contains 500 individuals from Wikipedia with varying levels of popularity, for which models are tasked to generate biographies; LongFact (Wei et al., 2024) extends Bios and includes 1,140 questions covering 38 manually-selected topics; WildHallu (Zhao et al., 2024) includes 7,917 entities derived from one million user-chatbot interactions in real-world settings.
Atomic Facts Generation and Verification. For all three datasets, we apply a FActScore-based (Min et al., 2023) factuality assessment approach. We first use GPT-4o to decompose the entire response into atomic facts, with each fact containing only a single piece of information. These atomic facts are then verified using GPT-4o, cross-referenced with evidence from Wikipedia and Google Search. The detailed prompts for generating atomic facts are provided in Appendix B.
Confidence Elicitation. We use P(true) (Kadavath et al., 2022), Self-Rating (Tian et al., 2023), and Semantic Entropy (SE) (Kuhn et al., 2022) as the baseline confidence elicitation methods. They are all calculated in response-level. For Gen-Binary, we apply the Llama3-8B-Instruct for better NLI performance. In WAvg, we assign a weighting 0.7 to generative methods and 0.3 to discriminative ones, based on validation results.
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Llama3-8B-Instruct
P(true) 30.2 45.1 46.0 18.9 16.3 16.8 40.5 25.7 26.1
Self-Rating 40.5 38.7 39.3 21.5 14.1 14.3 50.2 18.6 19.0
SE 42.1 37.4 38.0 23.0 13.5 13.8 52.0 17.8 18.2
[…]
Dis-Rating 73.8 25.7 25.7 34.1 2.9 2.8 71.7 3.6 3.8
Dis-Single 58.0 27.2 27.2 20.9 8.7 8.4 55.9 14.4 14.1
Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.8 7.2
Gen-Multi 71.4 35.8 36.0 37.5 11.6 11.8 62.6 22.0 24.1
Mistral-7B-Instruct
P(true) 32.8 44.5 44.9 22.0 16.7 17.0 41.2 24.3 25.2
Self-Rating 42.3 37.1 37.5 26.1 14.5 14.7 52.0 18.1 19.3
SE 44.1 36.5 36.7 28.0 13.8 14.1 53.4 17.4 18.5
[…]
Dis-Rating 55.0 41.4 41.5 40.8 4.4 4.2 60.4 16.9 16.7
Dis-Single 70.3 16.0 16.0 32.8 8.8 9.0 65.3 10.4 9.7
Gen-Binary 74.9 8.5 8.0 64.1 2.5 2.5 73.9 10.3 10.3
Gen-Multi 60.7 38.7 38.7 49.6 11.9 12.6 65.6 26.2 27.5
Qwen2-7B-Instruct
P(true) 33.5 45.0 44.8 28.3 11.2 11.6 35.4 12.7 12.9
Self-Rating 48.2 24.3 23.9 36.7 6.9 7.0 48.0 9.8 9.5
SE 49.8 22.9 22.6 38.9 6.5 6.4 49.2 8.9 9.2
[…]
Dis-Rating 63.0 40.7 41.0 29.9 4.4 4.4 54.0 9.1 8.9
Dis-Single 52.8 19.8 19.3 30.9 4.9 5.1 60.4 5.3 5.6
Gen-Binary 72.4 5.4 5.7 67.6 2.0 1.8 72.2 6.5 6.4
Gen-Multi 43.1 38.8 38.3 52.1 11.4 11.8 63.2 21.6 23.9
Table 2: Response Calibration Results. All the numbers are in percentages.
5.2 Results
Overall, the tested LLMs are not well-calibrated at the atomic fact level. Table 1 lists our main atomic calibration results, which indicate that different confidence elicitation methods yield varying calibration scores. Although there is no universally accepted threshold for low ECE, a well-calibrated model typically achieves an ECE close to 1%, as shown in (Guo et al., 2017) and (Zhu et al., 2023). However, even with the most robust method, Gen-Binary, the ECE scores remain around 10%, indicating a significant calibration gap. Among the models, Qwen2-7B-Instruct demonstrates slightly better calibration compared to the other two.
Atomic calibration can enhance macro calibration. Table 2 shows the main results of response-level calibration. For the five atomic-level methods, we calculate the average confidence of the facts in a response to obtain the response-level confidence. The results indicate that atomic calibration leads to better overall results compared to the baseline methods, highlighting the helpfulness of more fine-grained calibration analysis.
Confidence fusion can further improve both atomic and macro calibration. Table 4 presents the results of various confidence fusion strategies at the atomic level (with response-level results in Appendix D). The best results are consistently achieved by fusion methods. Among these strategies, the weighted average (WAvg) is the most effective. The benefits of the fusion strategies are further discussed in §6.1. Interestingly, we find that combining methods within the same confidence type (such as Dis-Rating with Dis-Context) does not lead to improved calibration (in Appendix D). A case study showing the effectiveness of confidence fusion is in Figure 4.
Larger model size does not necessarily result in better calibration. Table 3 compares the calibration levels of models with different sizes. Our two key findings are: (1) With generative methods, there is little difference in calibration between larger and smaller models; (2) With discriminative methods, larger models generally provide better calibration. We hypothesize that this is because discriminative methods require models to self-assess the confidence of their own outputs, and larger models typically possess stronger discriminative abilities (Saunders et al., 2022).
Bios LongFact WildHallu
ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
Gen-Binary
Llama3-8B-Instruct 10.0 17.8 83.1 8.5 11.4 77.3 11.1 15.2 82.0
Llama3-70B-Instruct 10.0 16.5 82.5 8.3 9.3 73.7 9.5 12.3 78.3
[…]
Mistral-8x7B-Instruct 12.3 18.5 79.8 7.8 9.0 76.3 9.8 13.4 77.8
[…]
Qwen2-57B-Instruct 10.5 18.1 82.3 7.8 10.0 78.3 9.2 13.6 81.7
Qwen2-72B-Instruct 11.2 16.6 83.4 7.6 8.3 76.6 8.6 11.9 77.7
Dis-Rating
Llama3-8B-Instruct 26.8 29.0 71.1 3.5 12.0 66.9 5.3 15.2 79.8
Llama3-70B-Instruct 10.6 19.3 73.2 4.2 8.0 74.2 4.3 11.5 81.2
[…]
Mistral-8x7B-Instruct 15.3 22.6 70.8 5.3 8.6 72.6 7.6 14.7 72.9
[…]
Qwen2-57B-Instruct 23.2 27.0 69.3 2.2 9.8 71.3 5.2 15.2 77.2
Qwen2-72B-Instruct 11.4 21.0 71.6 6.1 7.7 77.1 4.0 11.7 79.2
Table 3: Atomic Calibration Results. All the numbers are in percentages. Our two key findings are: (1) for generative methods, there is no significant difference in calibration between larger and smaller models; (2) however, with discriminative methods, larger models tend to provide better calibration.
Bios LongFact WildHallu
ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
Gen-Binary 10.0 17.8 83.1 8.5 11.4 77.3 11.1 15.2 82.0
Dis-Rating 26.8 29.0 71.1 3.5 12.0 66.9 5.3 15.2 79.8
Dis-Context 35.5 35.8 74.5 11.9 13.6 74.4 12.5 16.5 83.5
[…]
HMean 9.8 17.4 84.0 4.1 11.0 79.6 5.6 13.3 87.0
ProdConf 7.4 16.7 84.1 13.4 12.8 79.6 11.5 14.1 87.0
WAvg 10.9 17.4 84.4 3.3 10.3 79.9 5.1 13.0 87.0
Table 4: Atomic calibration results of confidence fusion strategies for Llama3-8B-Instruct, demonstrating the effectiveness of confidence fusion in combining generative and discriminative confidence estimates.
6 Discussion
6.1 Confidence Methods Alignment
(a) Atomic Level Confidences.
(b) Response Level Confidences.
Figure 2: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods in Llama3-8B-Instruct on WildHallu. Warmer colors are for higher correlations.
Figure 3: Average confidence scores across different parts of long-form responses. For discriminative methods, confidence decreases as the generation progresses, while generative methods show the lowest confidence in the middle sections.
To further explore the reasons behind the improvements provided by confidence fusion, we show the correlation between different confidence elicitation methods in Figure 2 (using WildHallu as the study case and more results are in Appendix E). Our findings are summarized as follows:
Confidence methods within the same type are better aligned. In Figure 2, warmer colors indicate higher Spearman correlation scores. Confidence elicitaton methods of the same type (top left for generative and bottom right for discriminative) show stronger correlations compared to those across different types. This helps to explain why fusing generative and discriminative methods are effective, since these two types capture different aspects of uncertainty and are complementary to each other.
The alignment is stronger at the response level than at the atomic level. When comparing atomic and macro calibration, we observe that the alignment is stronger for the latter. In atomic calibration, several methods display weak correlations (indicated in blue), while the correlations are generally higher at response level (indicated in red). Similarly, methods from different types show more disagreement than those of the same type. This highlights the need for future research on the discrepancies between generative and discriminative confidence elicitation methods, as well as how to better unify these approaches.
6.2 Confidence Across Different Positions
As each long-form response contains multiple atomic facts, we analyze how confidence and factuality scores evolve during the generation process. Specifically, we divide all atomic facts $C$ into five equal parts along the generation process. Part 1 represents the beginning of the generation, and part 5 corresponds to the end. We calculate the average confidence score for each part of the responses and present the results in Figure 3.
With discriminative methods, models exhibit decreasing confidence in atomic facts as the generation progresses. We observe similar trends across all discriminative methods. This contrasts with previous findings, which used logits as a measure of confidence and found that models tend to become more confident during long generation sequences (Zhang et al., 2023a). Our results show that discriminative methods indicate lower confidence in the model’s output toward the latter parts of the generation.
With generative methods, the model shows the lowest average confidence in the middle part of the generation.
We hypothesize that this is because the tested models tend to provide general introductions and conclusions at the beginning and the end of the generation. During consistency checking, these statements are frequently cross-referenced, leading to higher confidence. For example, in Bios, statements like “[a person] is famous” or “[a person] made a significant impact in his field” are often repeated across samples. On the contrary, in the middle parts where the models address more specific facts about individuals’ lives, careers and achievements, they tend to cover different aspects and details.
7 Related Work
Figure 4: An example from WildHallu dataset by Mistral-7B-Instruct. We only select five atomic facts for demonstration. The example shows the effectiveness of calculating confidence in atomic level with fusion strategy.
Atomic Facts Generation and Verification. Long-form responses often contain both correct and incorrect statements, which impact the overall factuality assessments. Min et al. (2023) propose breaking long responses into atomic facts and calculating the precision of these fact pieces to determine the overall factuality score. Wei et al. (2024) and Zhao et al. (2024) extend this paradigm by expanding the dataset to include more domains beyond biographies. Song et al. (2024) design VeriScore for diverse long-form generation tasks that feature both verifiable and unverifiable content. Chiang and Lee (2024) introduce D-FActScore, specifically designed for content with ambiguous entities. Overall, the approach of breaking long-form responses into atomic facts for verification is nowadays widely accepted in factuality assessment.
Uncertainty and Calibration in Long-form Generations. Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024). However, limited work has explored calibration for long-form generations. Huang et al. (2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses. Band et al. (2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation. Zhang et al. (2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs. None of the aforementioned studies have systematically examined calibration at a fine-grained level. Our work aims to fill this gap with a focus on factuality.
8 Conclusion
We introduce atomic calibration, a fine-grained approach for evaluating LLM calibration at the atomic claim level, addressing the limitations of traditional response-level calibration in long-form generation. Our experiments show that atomic calibration is well-suited for long-form generation and complements macro calibration. We also demonstrated that combining different confidence elicitation methods leads to better calibration results and atomic calibration can provide deeper insights into model confidence during the generation process. This approach offers a more reliable way to assess and improve the trustworthiness of LLM outputs in real-world applications.
Limitation
First, our work primarily focuses on the factuality aspect of LLMs. As mentioned in Section 2, the task $t$ can be various aspects of the quality of a long-form response, such as coherence, creativity, writing style, and more. Unlike previous studies that use the overall quality of long-form responses to evaluate calibration (Huang et al., 2024), we concentrate specifically on factuality in this paper. We argue that the hallucination problem is among the most significant challenges faced by LLMs (Zhang et al., 2023b; Huang et al., 2023).
Second, we test the calibration only on open-source LLMs for two main reasons: (1) After assessing the atomic and macro calibration levels of LLMs, our next step is to adjust the model to better reflect its confidence (i.e., for better calibration). Closed-source models are not directly applicable to this calibration process. (2) Our discrimination methods typically require logit access, which is generally unavailable in closed-source models. If logits are accessible, our methods can be directly applied to closed-source models without affecting the atomic calibration process.
Ethics Statement
Our research adheres to strict ethical standards. We ensured compliance with the licenses of all datasets and models used. No human participants were involved in our experiments. After thorough assessment, we do not anticipate any additional ethical concerns or risks related to our work.
Appendix
Appendix A Atomic Calibration Metrics
ECE
In computing the Expected Calibration Error (ECE), the predictions are sorted and divided into a fixed number of bins $K$. The predicted value of each test instance falls into one of the bins. $ECE$ uses empirical estimates as follows:
$ECE=\sum_{i=1}^{K}P(i)\cdot\left|o_{i}-e_{i}\right|,$
where $o_{i}$ is the true fraction of positive instances in bin $i$, $e_{i}$ is the mean of the post-calibrated probabilities for the instances in bin $i$, and $P(i)$ is the empirical probability (fraction) of all instances that fall into bin $i$. The lower the $ECE$ value is, the better a model is calibrated.
Brier Score
The Brier score measures the accuracy of probabilistic predictions. For unidimensional predictions, it is strictly equivalent to the mean squared error as applied to predicted probabilities. Suppose that on each of the $n$ occasions an event can occur in only one of $r$ possible classes or categories and on one such occasion $i$, the forecast probabilities are $f_{i1},f_{i2}$, $\ldots f_{ir}$, that the event will occur in classes $1,2,\ldots r$, respectively. The $r$ classes are chosen to be mutually exclusive and exhaustive so that
$\sum_{i=1}^{r}f_{ij}=1,i=1,2,3,\ldots n.$
The Brier Score $P$ is then defined by
$P=\frac{1}{n}\sum_{j=1}^{r}\sum_{i=1}^{n}\left(f_{ij}-E_{ij}\right)^{2},$
where $E_{ij}$ takes the value 1 or 0 according to whether the event occurred in class $j$ or not.
AUROC
Following (Kuhn et al., 2022), AUROC metric is equivalent to the probability that a randomly chosen correct answer has a higher confidence score than a randomly chosen incorrect answer. Higher scores are better for AUROC, and perfect confidence score is 1, while a random confidence measure would be 0.5.
Appendix B Prompts
Prompts to Generate Responses Bios: Tell me a bio of <ENTITY>. LongFact: Can you detail the concept of Gravitational Lensing and how it is utilized in modern astronomy for the study and understanding of the universe? Can you explain the process and importance of customer journey mapping in creating effective marketing strategies? WildHallu: In a paragraph, could you tell me what you know about <ENTITY>?
Table 5: Prompts to generate responses. We use the default prompt template for Bios and WildHallu. For LongFact, it has corresponding Prompt for each question, so we only list two examples here.
Gen-Binary Sample: {sample} Sentence: {sentence} Is the sentence supported by the sample above? You should answer the question purely based on the given sample and not your own knowledge. Do not output the explanations. Your answer should be within ‘‘yes’’ or ‘‘no’’. Answer:
Table 6: Prompt for Gen-Binary.
Gen-Multi Sample: {sample} Sentence: {sentence} Is the sentence supported, refuted, or not mentioned by the sample above? You should answer the question purely based on the given sample and not your own knowledge. Do not output the explanations. Your answer should be within ‘‘supported’’, ‘‘refuted’’, or ‘‘not mentioned’’. Answer:
Table 7: Prompt for Gen-Multi.
Dis-Single Please evaluate the truthfulness of the following statement based on your internal knowledge. Respond only with ‘‘True’’ or ‘‘False’’ with no other information. Statement: {atomic_fact_piece} True or False:
Table 8: Prompt for Dis-Single.
Dis-Context You are given a topic and a statement related to that topic. Evaluate the truthfulness of the statement based solely on your internal knowledge. A statement is true only if it is factually accurate according to what you know. Do not use the topic to evaluate the statement. Respond with ‘‘True’’ or ‘‘False’’ only, without any additional information. Topic: {context} Statement: {atomic_fact_piece} True or False:
Table 9: Prompt for Dis-Context.
Dis-Rating Please rate your confidence in the truth of the following statement on a scale from 0 to 10. A confidence of 0 means you are not confident at all whether the statement is true, while 10 means you are absolutely confident the statement is true. The topic provides context for the statement, but focus on the statement itself when rating your confidence. Provide only the numerical value for your confidence level, bounded by dollar signs (\$), without any additional information. Topic: {context} Statement: {atomic_fact_piece} Confidence:
Table 10: Prompt for Dis-Rating.
Appendix C Experiment Details
We use vLLM (Kwon et al., 2023) for our LLM inference tasks, with the following parameters: temperature = 1, top-$p$ = 0.95, and a maximum output of 512 tokens. For discriminative confidence elicitation methods, we set the temperature to 0 and only consider the top 10 logits. For generative methods, we use $N=20$ samples. The experiments are conducted on A100-SXM-40GB GPUs. Running the discriminative methods takes 30 minutes for 500 samples, while the generative methods take 1.3 hours for the same number of samples. We use GPT-4o as the auxiliary model for generating atomic claims and fact-checking the LLM.
Appendix D Confidence Fusion Results
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.2 7.8
Dis-Rating 73.8 25.7 25.7 34.1 2.8 2.9 71.7 3.8 3.6
Dis-Context 55.4 33.2 34.0 29.7 5.4 5.6 65.9 9.1 9.5
[…]
HMean 81.7 11.5 11.5 54.5 2.0 2.0 78.7 3.2 3.0
ProdConf 82.2 7.7 7.1 54.9 14.2 14.3 78.7 12.2 12.1
WAvg 81.8 10.9 11.0 55.6 1.5 0.9 77.2 5.0 5.0
Table 11: Macro calibration results of confidence fusion strategies for Llama3-8B-Instruct.
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 74.9 8.0 8.5 64.1 2.5 2.5 73.9 10.3 10.3
Dis-Rating 55.0 41.5 41.4 40.8 4.2 4.4 60.4 16.7 16.9
Dis-Context 79.7 8.2 8.3 47.9 4.2 4.1 72.3 5.7 6.1
[…]
HMean 83.5 5.8 6.2 62.5 1.9 2.1 77.9 6.5 6.6
ProdConf 83.8 13.1 13.0 62.5 10.6 10.4 78.0 7.2 7.1
WAvg 82.0 8.0 8.0 65.1 1.8 2.0 78.4 8.9 8.9
Table 12: Macro calibration results of confidence fusion strategies for Mistral-7B-Instruct.
Bios LongFact WildHallu
SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
Gen-Binary 72.4 5.7 5.4 67.6 1.8 2.0 72.2 6.4 6.5
Dis-Rating 63.0 41.0 40.7 29.9 4.4 4.4 54.0 8.9 9.1
Dis-Context 66.5 14.4 14.8 40.6 3.9 3.9 66.8 3.7 4.0
[…]
HMean 76.2 7.8 8.0 62.4 1.4 1.4 74.7 2.3 2.2
ProdConf 76.0 12.9 12.6 62.5 10.4 10.4 74.9 12.7 12.7
WAvg 75.7 7.5 8.4 66.4 1.2 1.4 75.6 4.5 4.3
Table 13: Macro calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
Bios LongFact WildHallu
ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
Gen-Binary 13.7 19.0 81.9 8.4 11.5 80.1 12.7 17.0 81.3
Dis-Rating 44.5 42.5 65.0 10.0 14.2 67.9 19.7 23.9 68.1
Dis-Context 24.8 26.0 77.5 15.7 16.1 75.3 20.6 21.7 79.8
[…]
HMean 14.3 18.3 82.1 7.6 12.2 81.0 11.5 16.6 83.4
ProdConf 14.2 18.3 82.3 9.5 13.0 81.0 7.9 15.9 83.5
WAvg 10.6 16.6 84.7 5.5 10.7 82.1 12.3 16.4 84.4
Table 14: Atomic calibration results of confidence fusion strategies for Mistral-7B-Instruct.
Bios LongFact WildHallu
ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
Gen-Binary 10.9 16.7 83.8 6.3 9.9 81.9 9.5 14.0 82.5
Dis-Rating 41.5 39.7 64.2 3.5 11.7 62.6 8.2 18.1 70.4
Dis-Context 26.5 28.3 75.5 13.9 14.8 77.9 17.2 19.4 81.2
[…]
HMean 11.1 16.9 82.7 2.7 9.6 81.7 4.9 13.6 83.8
ProdConf 12.4 17.0 83.1 8.3 10.6 81.7 7.2 13.7 83.9
WAvg 10.7 15.9 84.8 2.6 9.2 82.8 6.8 13.5 84.3
Table 15: Atomic calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
Appendix E Confidence Alignment
(a) Mistral-7B-Instruct : atomic
(b) Mistral-7B-Instruct : response
(c) Qwen2-7B-Instruct : atomic
(d) Qwen2-7B-Instruct : response
(e) Llama3-8B-Instruct : atomic
(f) Llama3-8B-Instruct : response
Figure 5: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for Bios. Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
(a) Mistral-7B-Instruct : atomic
(b) Mistral-7B-Instruct : response
(c) Qwen2-7B-Instruct : atomic
(d) Qwen2-7B-Instruct : response
(e) Llama3-8B-Instruct : atomic
(f) Llama3-8B-Instruct : response
Figure 6: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for LongFact. Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
(a) Mistral-7B-Instruct : atomic
(b) Mistral-7B-Instruct : response
(c) Qwen2-7B-Instruct : atomic
(d) Qwen2-7B-Instruct : response
(e) Llama3-8B-Instruct : atomic
(f) Llama3-8B-Instruct : response
Figure 7: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for WildHallu. Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.

### v2

@@
 Large language models (LLMs) often suffer from hallucinations, posing significant challenges for real-world applications.
-Confidence calibration, which estimates the underlying uncertainty of model predictions, is essential to enhance the LLMs’ trustworthiness.
-Existing research on LLM calibration has primarily focused on short-form tasks, providing a single confidence score at the response level (macro calibration).
-However, this approach is insufficient for long-form generations, where responses often contain more complex statements and may include both accurate and inaccurate information.
-Therefore, we introduce atomic calibration, a novel approach that evaluates factuality calibration at a fine-grained level by breaking down long responses into atomic claims.
-We classify confidence elicitation methods into discriminative and generative types and demonstrate that their combination can enhance calibration.
-Our extensive experiments on various LLMs and datasets show that atomic calibration is well-suited for long-form generation and can also improve macro calibration results.
-Additionally, atomic calibration reveals insightful patterns in LLM confidence throughout the generation process.
+Confidence calibration, as an effective indicator of hallucination, is thus essential to enhance the trustworthiness of LLMs.
+Prior work mainly focuses on short-form tasks using a single response-level score (macro calibration), which is insufficient for long-form outputs that may contain both accurate and inaccurate claims.
+In this work, we systematically study atomic calibration, which evaluates factuality calibration at a fine-grained level by decomposing long responses into atomic claims.
+We further categorize existing confidence elicitation methods into discriminative and generative types, and propose two new confidence fusion strategies to improve calibration.
+Our experiments demonstrate that LLMs exhibit poorer calibration at the atomic level during long-form generation.
+More importantly, atomic calibration uncovers insightful patterns regarding the alignment of confidence methods and the changes of confidence throughout generation.
+This sheds light on future research directions for confidence estimation in long-form generation.
 Atomic Calibration of LLMs in Long-Form Generations
@@
 We only list three atomic claims for illustration.
-While large language models (LLMs), such as Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023), and GPT models (OpenAI, 2022), excel in various tasks, they still struggle with faithfulness and reliability issues.
+While large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023; OpenAI, 2022) excel in various tasks, they still struggle with trustworthiness issues.
 LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023).
@@
 A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024).
-Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024).
-These studies often use datasets like TriviaQA (Joshi et al., 2017) and Natural Questions (Joshi et al., 2017), where answers typically have fewer than 10 words.
-However, in real-world applications, responses to user queries are often much longer (Zhang et al., 2024), sometimes extending to hundreds or even thousands of words.
-In such cases, the quality of LLM responses is not simply binary (correct or incorrect), as answers may include both accurate and inaccurate statements.
-Previous work on long-form calibration has primarily focused on response-level calibration (Zhang et al., 2024; Huang et al., 2024) (which we term macro calibration).
-This approach provides an overall confidence estimation for the entire response (as shown in the upper part of Figure 1).
-However, it is insufficient for long-form generation, as it cannot adequately capture the model’s fine-grained uncertainty across multiple factual statements.
-In this work, we introduce the concept of atomic calibration, which evaluates calibration at the fine-grained level of atomic claims (as illustrated in the lower part of Figure 1).
-We focus on the aspect of factuality, since hallucination are a widely recognized issue of LLMs (Zhang et al., 2023b; Huang et al., 2023) and the factuality of statements can be objectively determined.
-Accurate factual calibration is crucial for mitigating the hallucination problem (Mahaut et al., 2024).
-To perform atomic calibration, we decompose long responses into self-contained atomic claims and evaluate the calibration for all atomic claims.
-Various confidence elicitation methods are studied to provide accurate confidence estimation.
-We provide a detailed investigation into long-form generation calibration, covering 7 LLMs and 3 datasets.
-The main contributions of our work are as follows:
-  • We propose the novel concept of atomic calibration.
-Unlike macro calibration, which provides a single response-level confidence score, atomic calibration offers a more fine-grained analysis of model calibration, making it better suited for long-form generation (§2).
-  • We introduce a categorization of confidence elicitation methods for long-form generation into two types: discriminative and generative, estimating the model’s intrinsic and external confidence, respectively.
-We examine five confidence elicitation methods and assess their effectiveness for atomic calibration (§3).
-  • We show that atomic calibration is particularly well-suited for long-form generation, and that macro calibration can be enhanced by incorporating atomic calibration results.
-We also find the two types of confidence elicitation methods are complementary and their fusion can provide better calibration results.
-Atomic calibration also enables deeper analysis, revealing patterns in confidence and calibration changes throughout the generation process (§5).
+Improved calibration enables more reliable confidence estimation, warning users when not to trust the model and thus mitigating the impact of hallucinations.
+Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024), using datasets like TriviaQA and Natural Questions (Joshi et al., 2017), where answers are typically under 10 words.
+In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024; Yang et al., 2024b), spanning hundreds or thousands of words.
+In such cases, response quality is not simply binary, as answers may mix accurate and inaccurate statements.
+Recent work has begun to address calibration in long-form generation (Zhang et al., 2024; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024).
+Several approaches estimate a single confidence score for the entire response (macro calibration; upper, Figure 1), while others assess confidence at the level of atomic claims (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024) (atomic calibration; lower, Figure 1).
+However, previous work leaves several key research questions unanswered: Why is it important to evaluate calibration at the atomic-claim level?
+What factors influence calibration results at this level?
+What patterns can be observed by analyzing calibration at the atomic-claim level?
+In this work, we systematically examine atomic calibration: A long response is decomposed into atomic claims, each containing a single factual statement, and confidence scores are assigned using various elicitation methods.
+To analyze different confidence elicitation methods, we categorize them into discriminative (intrinsic confidence estimation) and generative (external confidence assessment).
+Our experiments on three long-form QA datasets with seven LLMs reveal that: (1) Models that appear well-calibrated at the response level perform poorly at the atomic level (Figure 2, Table 1); (2) Leveraging atomic calibration enhances macro calibration (Table 2).
+These two reasons highlight the need for research on atomic-level calibration to develop better-calibrated models.
+We further investigate the characteristics of discriminative and generative confidence.
+Our analysis yields two main findings: (1) Discriminative and generative methods are complementary; combining them improves calibration, while combinations within the same category offer limited gains.
+(2) Generative methods maintain consistent calibration across different model sizes, whereas discriminative methods benefit from increased model sizes.
+Motivated by finding (1), we propose two novel fusion strategies based on confidence agreement to integrate generative and discriminative confidence.
+Our strategies outperform existing fusion methods.
+Our atomic-level analysis (Section 6) offers deeper insights into confidence method alignment and confidence changes during generation.
+Confidence methods within the same category align better, explaining why cross-category fusion is more effective.
+Interestingly, with discriminative methods, model confidence in atomic facts tends to decrease as generation progresses.
+In contrast, generative methods show the lowest average confidence in the middle of the generation process.
+These results highlight the necessity of fine-grained calibration evaluation for long-form generation, given us insights on model trustworthiness and usability.
+2 Related Work
+Atomic Claims Generation and Verification.
+Long-form responses often contain both correct and incorrect statements, which impact the overall factuality assessments.
+Min et al.
+(2023) propose breaking long responses into atomic facts and calculating the precision of these fact pieces to determine the overall factuality score.
+Wei et al.
+(2024) and Zhao et al.
+(2024) extend this paradigm by expanding the dataset to include more domains beyond biographies.
+Song et al.
+(2024) design VeriScore for diverse long-form generation tasks that feature both verifiable and unverifiable content.
+Chiang and Lee (2024) introduce D-FActScore, specifically designed for content with ambiguous entities.
+Decomposing long-form responses into atomic claims and fact-checking them individually has become a widely adopted pipeline.
+Uncertainty and Calibration in Long-form Generations.
+Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024).
+There is an increasing interest on calibration for long-form generations.
+Huang et al.
+(2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses.
+Band et al.
+(2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation.
+Zhang et al.
+(2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs.
+Another line of work (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024; Yuan et al., 2024) decomposes sentences into atomic claims and assigns confidence scores to each claim.
+However, a unified definition of atomic-level calibration remains lacking.
+Clarifying this concept and identifying key influencing factors are essential steps toward improving calibration in long-form generation.
 3 Atomic Calibration
-For a language model $\mathcal{M}$, let $x\in\mathcal{X}$ represent the response generated by model $\mathcal{M}$ for a query $q$, and let $y\in\mathcal{Y}_{t}$ denote the corresponding label, where $\mathcal{Y}_{t}\subseteq[0,1]$ indicates a quality score for a specific task $t\in T$.
-Note that unlike multiple-choice and short-form questions, which typically only focus on the correctness of the answer, the tasks in $T$ involve various aspects, such as factuality, coherence, creativity, etc.
-We define a probability prediction function $f:\mathcal{X}\to\Delta^{\lvert\mathcal{Y}_{t}\rvert}$, where $\Delta^{\lvert\mathcal{Y}_{t}\rvert}$ represents the $\lvert\mathcal{Y}_{t}\rvert$-dimensional simplex.
-In this context, $f(x)_{y}$ denotes the probability assigned to the label $y$ given the generated output $x$.
-In the remaining of this work, we focus specifically on calibrating for factuality.
-In this case, $\mathcal{Y}$ represent $\mathcal{Y}_{t}$ when $t$ corresponds to factuality.
-$\mathcal{Y}$ takes values in [0, 1], indicating the factuality level of a response.
-Based on the calibration concepts discussed in (Guo et al., 2017), we define the calibration of each response as follows:
+For a language model $\mathcal{M}$, let $x\sim M(x\mid q)$ denote the response generated by $\mathcal{M}$ for a query $q$, $x\in\mathcal{X}$.
+Let $y\in\mathcal{Y}_{t}$ be the corresponding label, representing a quality score ranging from 0 to 1 for a specific task $t\in T$.
+Unlike multiple-choice or short-form questions, which mainly assess correctness, tasks in $T$ cover diverse dimensions such as factuality, coherence, and creativity.
+We define a probability prediction function $f:\mathcal{X}\to\Delta^{\lvert\mathcal{Y}_{t}\rvert}$, where $\Delta^{\lvert\mathcal{Y}_{t}\rvert}$ denotes the $\lvert\mathcal{Y}_{t}\rvert$-dimensional probability simplex.
+Here, $f(x)_{y}$ represents the probability assigned to label $y$ for a generated output $x$.
+In this work, we focus on calibrating factuality, as hallucinations are a well-known issue in LLMs (Zhang et al., 2023b; Huang et al., 2023), and the factuality of atomic claims can be assessed objectively.
+In this setting, $\mathcal{Y}$ denotes $\mathcal{Y}_{t}$ for the factuality task $t$, where $\mathcal{Y}\subseteq[0,1]$ reflects the factuality level of a response.
+Following Guo et al.
+(2017), we define the calibration of each response as follows:
 Definition 1 (Macro Calibration on Factuality)
 A language model $\mathcal{M}$ that produces generations $x\sim\mathcal{M}(x\mid q)$ is said to be response-level (macro) calibrated if
-$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta,\quad\forall\beta\in\Delta^{\lvert\mathcal{Y}\rvert}.$
+$\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta,\quad\forall\beta\in[0,1].$
 In the context of long-form generation, a single response $x$ may encompass multiple atomic claims.
@@
 A language model $\mathcal{M}$, which generates a long-form response $x$ conditioned on the query $q$, $x\sim\mathcal{M}(x\mid q)$, is considered atomic-level calibrated if, for each atomic claim $c_{i}$ with its corresponding label $y_{i}$, the following condition holds:
-$\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta_{i},\quad\forall\beta_{i}\in\Delta^{\lvert\mathcal{Y}_{i}\rvert}.$
+$\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta_{i},\quad\forall\beta_{i}\in[0,1].$
 Remarks:
@@
 4 Confidence Elicitation Methods
-To calculate the confidence scores for model outputs in long-form generation, we define two types of confidence elicitation methods: generative and discriminative.
-Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty.
-The more frequently certain facts are covered by the samples, the higher the model’s confidence in those facts.
-In contrast, discriminative methods assess uncertainties by asking the model itself.
-This is motivated by the findings that models tend to perform better on discriminative tasks (Saunders et al., 2022), and thus they may already possess the capability to estimate the confidence of their own outputs in a discriminative manner.
-We first generate one response $x$ as the answer to the query $q$.
-Then, $x$ is broken into atomic claims $C$.
-Following (Min et al., 2023; Wei et al., 2024; Zhao et al., 2024), each atomic claim contains a single piece of information and must be self-contained.
+In this section, we define two types of confidence elicitation methods: generative and discriminative.
+We then introduce two novel confidence fusion strategies that considers confidence agreement when combining confidence scores.
+For the response $x$ to a query $q$, $x$ is broken into atomic claims $C$.
+Following previous work (Min et al., 2023; Wei et al., 2024; Zhao et al., 2024), each atomic claim contains a single piece of information and must be self-contained.
 For generative methods, we sample an additional set of responses $K$, and compare them against the original response $x$.
@@
 4.1 Generative Methods
+Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024; Jiang et al., 2024).
+Generally, an additional natural language inference (NLI) model is used to calculate the consistency.
+In particular, we have the following two variations:
 Gen-Binary.
-The basic assumption here is that if a fact is frequently conveyed when sampled multiple times, the model is considered “confident” about that fact.
-For an atomic claim $c_{i}$ in $C$, we utilize a natural language inference model $\mathcal{M}_{\mathrm{NLI}}$ to examine whether $c_{i}$ is supported or not supported by each of the additional samples.
+The basic assumption is that if a fact is frequently conveyed when sampled multiple times, the model is considered “confident” about that fact.
+For an atomic claim $c_{i}$ in $C$, we utilize a NLI model $\mathcal{M}_{\mathrm{NLI}}$ to examine whether $c_{i}$ is supported or not supported by each of the additional samples.
 Let $K_{s}$ be the set of samples supporting $c_{i}$.
-Then, the confidence in $c_{i}$ is calculated as:
-$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K\rvert}$
+Then, the confidence in $c_{i}$ is calculated as
+$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K\rvert}.$
 Gen-Multi.
 Gen-Multi assumes that the model is more confident in facts that are consistently expressed.
-Unlike Gen-Binary, it further divides the “not supported” class $K_{ns}$ into “conflict” ($K_{c}$) if the fact is presented differently in the sample, and “not mentioned” ($K_{nm}$) if the fact is not mentioned in the sample.
+Unlike Gen-Binary, it further divides the “not supported” ($K_{ns}$) into “conflict” ($K_{c}$) if the fact is presented differently in the sample, and “not mentioned” ($K_{nm}$) if the fact is not mentioned in the sample.
 We then calculate the confidence by only considering supporting and conflicting samples:
-$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K_{s}\rvert+\lvert K_{c}\rvert}$
+$Conf(c_{i},K)=\frac{\lvert K_{s}\rvert}{\lvert K_{s}\rvert+\lvert K_{c}\rvert}.$
 4.2 Discriminative Methods
+Discriminative methods assess uncertainties by asking the model itself (Tian et al., 2023; Xiong et al., 2023).
+This is motivated by the findings that models tend to perform better on discriminative tasks (Saunders et al., 2022), and thus they may already possess the capability to estimate the confidence of their own outputs in a discriminative manner.
 Dis-Single.
-Following (Kadavath et al., 2022; Tian et al., 2023), we directly ask the model whether one single atomic claim is true or false.
+Following Kadavath et al.
+(2022); Tian et al.
+(2023), we directly ask the model whether one single atomic claim is true or false.
 The probability the model assigns to token “True” ($P(true)$) in its generation is viewed as the confidence.
@@
 However, Tian et al.
-(2023) show that LLMs express uncertainty as effectively, or even more effectively, using numerical values rather than words.
+(2023) demonstrate LLMs achieve comparable or even better results using numerical values.
 4.3 Confidence Fusion Strategies
-Since the generative and discriminative methods provide different perspectives of confidence estimation, it is natural to explore fusion methods that combine different confidence estimates for better calibration.
-After exploration, we consider four simple but effective options: (1) MinConf selects the minimum of the two confidence values, providing a conservative estimate; (2) HMean calculates the harmonic mean, providing a balanced fusion that weights lower confidence values more heavily, to help prevent overconfident predictions; (3) ProdConf multiplies the confidences, ensuring a high final confidence only when both estimates are strong, thereby effectively penalizing discrepancies between the two estimates; and (4) WAvg computes the weighted average, with a hyper-parameter weight, prioritizing the more reliable estimate and allowing flexibility to adjust the influence of each confidence source based on prior knowledge or validation results.
-4 Evaluation Metrics
-4.1 Atomic Calibration
-In atomic calibration, we utilize FActScore (Min et al., 2023) to measure the atomic claims’ factuality.
-For each claim, we have its binary factuality label and a continuous confidence score.
-Since this corresponds to standard calibration scenarios, we follow the conventions from previous work (Kuhn et al., 2022; Zhu et al., 2023; Tian et al., 2023) and use Expected Calibration Error (ECE) (Naeini et al., 2015), Brier Score (Brier, 1950), and AUROC to assess calibration.
-Details of the three metrics are in Appendix A.
-4.2 Macro Calibration
-In macro calibration, factuality scores for responses are represented as continuous values from 0 to 1, therefore, the atomic calibration metrics mentioned above are not directly applicable.
-Apart from the Spearman Correlation, we propose two novel metrics UCCE and QCCE that extend ECE to continuous factuality scores.
-Spearman Correlation.
-Following (Zhang et al., 2024; Huang et al., 2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores.
-Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
-Uniform Continuous Calibration Error (UCCE).
-UCCE evaluates how well a model’s continuous predictions align with the true values by dividing the predicted values into bins of equal size.
-This metric assesses the model’s calibration across the range of predictions.
-The calibration error is computed by comparing the mean predicted and true values within each bin.
-Let $\hat{y}_{i}$ represent the predicted value for the $i$-th sample, $y_{i}$ the true value for the $i$-th sample, $B_{m}$ the set of samples falling into the $m$-th bin, $M$ the number of bins, and $N$ the total number of samples.
-The UCCE is computed as follows:
-$\text{UCCE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\frac{\sum_{i\in B_{m}}\hat{y}_{i}}{|B_{m}|}-\frac{\sum_{i\in B_{m}}y_{i}}{|B_{m}|}\right|.$
-UCCE is particularly effective for scenarios where predicted values are uniformly distributed.
-Quantile Continuous Calibration Error (QCCE).
-QCCE uses a similar approach to UCCE but utilizes quantile binning, ensuring that each bin contains an equal number of samples.
-Consequently, each bin has a size of $\lvert B_{m}\rvert=\frac{N}{M}$.
-QCCE is advantageous for skewed or non-uniform distributions of predicted values, as it avoids sparsity in any bin by ensuring a balanced representation across all bins.
-The QCCE is calculated as follows:
-QCCE $=\frac{1}{M}\sum_{m=1}^{M}\left|\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{y}_{i}-\frac{1}{|B_{m}|}\sum_{i\in B_{m}}y_{i}\right|$
-$=\frac{1}{N}\sum_{m=1}^{M}\left|\sum_{i\in B_{m}}\hat{y}_{i}-\sum_{i\in B_{m}}y_{i}\right|.$
+Combining confidence scores has proven effective for calibration (Huang et al., 2024; Rivera et al., 2024), but existing methods typically only use a single fixed weight, $\alpha$, to combine the scores, ignoring the confidence disagreement.
+For instance, Rivera et al.
+(2024) computes the weighted average (WAvg) for confidence scores $A$ and $B$: $C=A\cdot\alpha+B\cdot(1-\alpha)$.
+This approach does not account for the agreement between the two scores.
+For example, when $\alpha=0.5$, confidences of 0 and 1 are treated the same as confidences of 0.4 and 0.6, although the former may indicate higher uncertainty due to a larger disagreement.
+To address this, we propose two simple but effective methods that consider confidence disagreement $d=B-A$.
+AdjustedAlpha adjusts the weight $\alpha$ based on the confidence difference:
+$\alpha^{\prime}=\alpha+\gamma_{\text{a}}\cdot d,$
+where $\gamma_{\text{a}}$ is a small constant (e.g., 0.1), followed by $C^{\prime}=A\cdot\alpha^{\prime}+B\cdot(1-\alpha^{\prime})$.
+DampedFusion applies a damping factor based on the agreement:
+$\gamma(d)=1-k\cdot|d|,$
+where $k$ is a small constant (e.g., 0.02) that controls the damping sensitivity, followed by $C^{\prime}=C\cdot\gamma(d)$.
+For baselines, we also include: MinConf, which selects the minimum confidence; HMean, which calculates the harmonic mean; and ProdConf, which multiplies the confidences.
 Bios LongFact WildHallu
@@
 Models.
-We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024).
+We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024a).
 Datasets.
@@
 For all three datasets, we apply a FActScore-based (Min et al., 2023) factuality assessment approach.
-We first use GPT-4o to decompose the entire response into atomic facts, with each fact containing only a single piece of information.
+We first use GPT-4o to decompose the entire response into atomic facts.
 These atomic facts are then verified using GPT-4o, cross-referenced with evidence from Wikipedia and Google Search.
-The detailed prompts for generating atomic facts are provided in Appendix B.
+The detailed prompts for generating atomic facts are provided in Appendix I.
 Confidence Elicitation.
-We use P(true) (Kadavath et al., 2022), Self-Rating (Tian et al., 2023), and Semantic Entropy (SE) (Kuhn et al., 2022) as the baseline confidence elicitation methods.
+We use P(true) (Kadavath et al., 2022), Self-Rating (Tian et al., 2023), Semantic Entropy (SE) (Kuhn et al., 2022), and Sum of Eigenvalues (EigV) (Lin et al., 2023) as the baseline confidence elicitation methods.
 They are all calculated in response-level.
-For Gen-Binary, we apply the Llama3-8B-Instruct for better NLI performance.
-In WAvg, we assign a weighting 0.7 to generative methods and 0.3 to discriminative ones, based on validation results.
+For Gen-Binary, we apply the Llama-3-8B-Instruct for better NLI performance.
+For WAvg, AdjustedAlpha, and DampedConf, we use a separate validation set for hyper-parameter tuning.
+Metrics.
+We use Expected Calibration Error (ECE) (Naeini et al., 2015) and Brier Score (BS) (Brier, 1950) as the primary metrics.
+These metrics are applicable to both atomic and macro calibration (see details in Appendix A), enabling a direct comparison between them.
+Additionally, we include AUROC to evaluate atomic calibration and Spearman Correlation for a more instance-specific assessment in macro calibration.
+Figure 2: Comparison of atomic level and response-level calibration for ECE and Brier Score.
+Atomic-level performance is generally worse than response-level performance, with data points consistently lying above the identity line.
 Bios LongFact WildHallu
-SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
+ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ SC $\uparrow$
 Llama3-8B-Instruct
-P(true) 30.2 45.1 46.0 18.9 16.3 16.8 40.5 25.7 26.1
-Self-Rating 40.5 38.7 39.3 21.5 14.1 14.3 50.2 18.6 19.0
-SE 42.1 37.4 38.0 23.0 13.5 13.8 52.0 17.8 18.2
+P(true) 45.1 25.9 30.2 16.3 4.8 18.9 25.7 13.5 40.5
+Self-Rating 38.7 23.4 40.5 14.1 4.2 21.5 18.6 12.9 50.2
+SE 37.4 21.8 42.1 13.5 3.4 23.0 17.8 11.7 52.0
+EigV 36.8 21.2 43.0 13.0 3.2 23.8 17.2 11.3 53.0
 […]
-Dis-Rating 73.8 25.7 25.7 34.1 2.9 2.8 71.7 3.6 3.8
-Dis-Single 58.0 27.2 27.2 20.9 8.7 8.4 55.9 14.4 14.1
-Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.8 7.2
-Gen-Multi 71.4 35.8 36.0 37.5 11.6 11.8 62.6 22.0 24.1
+Dis-Rating 25.7 11.7 73.8 2.9 1.6 34.1 3.6 3.5 71.7
+Dis-Single 27.2 13.7 58.0 8.7 2.6 20.9 14.4 7.3 55.9
+Gen-Binary 5.6 3.3 79.8 3.0 1.1 52.7 7.8 4.6 70.0
+Gen-Multi 35.8 18.1 71.4 11.6 2.7 37.5 22.0 10.5 62.6
 Mistral-7B-Instruct
-P(true) 32.8 44.5 44.9 22.0 16.7 17.0 41.2 24.3 25.2
-Self-Rating 42.3 37.1 37.5 26.1 14.5 14.7 52.0 18.1 19.3
-SE 44.1 36.5 36.7 28.0 13.8 14.1 53.4 17.4 18.5
+P(true) 44.5 27.1 32.8 16.7 7.4 22.0 24.3 19.8 41.2
+Self-Rating 37.1 26.4 42.3 14.5 6.5 26.1 18.1 14.5 52.0
+SE 36.5 23.9 44.1 13.8 3.7 28.0 17.4 14.3 53.4
+EigV 35.9 23.3 45.3 13.3 3.5 36.5 16.9 13.9 54.4
 […]
-Dis-Rating 55.0 41.4 41.5 40.8 4.4 4.2 60.4 16.9 16.7
-Dis-Single 70.3 16.0 16.0 32.8 8.8 9.0 65.3 10.4 9.7
-Gen-Binary 74.9 8.5 8.0 64.1 2.5 2.5 73.9 10.3 10.3
-Gen-Multi 60.7 38.7 38.7 49.6 11.9 12.6 65.6 26.2 27.5
+Dis-Rating 41.4 22.7 55.0 4.4 1.7 40.8 16.9 9.8 60.4
+Dis-Single 16.0 6.6 70.3 8.8 3.1 32.8 10.4 6.5 65.3
+Gen-Binary 8.5 3.8 74.9 2.5 1.0 64.1 10.3 5.0 73.9
+Gen-Multi 38.7 20.1 60.7 11.9 2.8 49.6 26.2 13.4 65.6
 Qwen2-7B-Instruct
-P(true) 33.5 45.0 44.8 28.3 11.2 11.6 35.4 12.7 12.9
-Self-Rating 48.2 24.3 23.9 36.7 6.9 7.0 48.0 9.8 9.5
-SE 49.8 22.9 22.6 38.9 6.5 6.4 49.2 8.9 9.2
+P(true) 45.0 27.9 33.5 11.2 5.6 28.3 12.7 15.6 35.4
+Self-Rating 24.3 25.1 48.2 6.9 4.9 36.7 9.8 14.7 48.0
+SE 22.9 23.1 49.8 6.5 3.7 38.9 8.9 13.9 49.2
+EigV 22.4 22.5 50.7 6.2 3.5 39.8 8.5 13.5 50.2
 […]
-Dis-Rating 63.0 40.7 41.0 29.9 4.4 4.4 54.0 9.1 8.9
-Dis-Single 52.8 19.8 19.3 30.9 4.9 5.1 60.4 5.3 5.6
-Gen-Binary 72.4 5.4 5.7 67.6 2.0 1.8 72.2 6.5 6.4
-Gen-Multi 43.1 38.8 38.3 52.1 11.4 11.8 63.2 21.6 23.9
-Table 2: Response Calibration Results. All the numbers are in percentages.
+Dis-Rating 40.7 21.3 63.0 4.4 1.9 29.9 9.1 6.3 54.0
+Dis-Single 19.8 8.4 52.8 4.9 2.4 30.9 5.3 4.8 60.4
+Gen-Binary 5.4 3.2 72.4 2.0 0.9 67.6 6.5 3.2 72.2
+Gen-Multi 38.8 20.0 43.1 11.4 2.6 52.1 21.6 9.5 63.2
+Table 2: Macro Calibration Results. All the numbers are in percentages.
 5.2 Results
 Overall, the tested LLMs are not well-calibrated at the atomic fact level.
-Table 1 lists our main atomic calibration results, which indicate that different confidence elicitation methods yield varying calibration scores.
+Table 1 lists our main atomic calibration results.
 Although there is no universally accepted threshold for low ECE, a well-calibrated model typically achieves an ECE close to 1%, as shown in (Guo et al., 2017) and (Zhu et al., 2023).
@@
 Among the models, Qwen2-7B-Instruct demonstrates slightly better calibration compared to the other two.
+Models that appear well-calibrated at the response level still perform poorly at the atomic level
+Figure 2 compares atomic and response-level scores for ECE and Brier Score across different datasets and confidence types.
+The data points consistently lie above the identity line, indicating that atomic-level errors are higher than response-level errors.
+This suggests that atomic calibration is crucial for fine-grained evaluation.
 Atomic calibration can enhance macro calibration.
@@
 The results indicate that atomic calibration leads to better overall results compared to the baseline methods, highlighting the helpfulness of more fine-grained calibration analysis.
-Confidence fusion can further improve both atomic and macro calibration.
-Table 4 presents the results of various confidence fusion strategies at the atomic level (with response-level results in Appendix D).
-The best results are consistently achieved by fusion methods.
-Among these strategies, the weighted average (WAvg) is the most effective.
-The benefits of the fusion strategies are further discussed in §6.1.
-Interestingly, we find that combining methods within the same confidence type (such as Dis-Rating with Dis-Context) does not lead to improved calibration (in Appendix D).
-A case study showing the effectiveness of confidence fusion is in Figure 4.
+The confidence fusion method considering confidence agreement outperforms other methods.
+Table 4 presents the results of various confidence fusion strategies at the atomic level (more results in Appendix F).
+The best performance is consistently achieved by AdjustedAlpha and DampedFusion.
+Notably, we observe that combining methods of the same confidence type (e.g., Dis-Rating with Dis-Context) does not lead to improved calibration.
+A case study demonstrating the effectiveness of confidence fusion is shown in Figure 11.
 Larger model size does not necessarily result in better calibration.
@@
 Qwen2-72B-Instruct 11.4 21.0 71.6 6.1 7.7 77.1 4.0 11.7 79.2
-Table 3: Atomic Calibration Results.
-All the numbers are in percentages.
-Our two key findings are: (1) for generative methods, there is no significant difference in calibration between larger and smaller models; (2) however, with discriminative methods, larger models tend to provide better calibration.
+Table 3: Atomic calibration results with different model sizes. All the numbers are in percentages.
 Bios LongFact WildHallu
@@
 WAvg 10.9 17.4 84.4 3.3 10.3 79.9 5.1 13.0 87.0
-Table 4: Atomic calibration results of confidence fusion strategies for Llama3-8B-Instruct, demonstrating the effectiveness of confidence fusion in combining generative and discriminative confidence estimates.
+[…]
+DampedFusion 5.0 15.6 84.7 3.5 9.8 80.0 4.8 12.4 87.9
+Table 4: Atomic calibration results of different confidence fusion strategies for Llama3-8B-Instruct. The fusion results are based on Gen-Binary and Dis-Rating.
 6 Discussion
 6.1 Confidence Methods Alignment
-(a) Atomic Level Confidences.
-(b) Response Level Confidences.
-Figure 2: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods in Llama3-8B-Instruct on WildHallu.
-Warmer colors are for higher correlations.
-Figure 3: Average confidence scores across different parts of long-form responses.
-For discriminative methods, confidence decreases as the generation progresses, while generative methods show the lowest confidence in the middle sections.
-To further explore the reasons behind the improvements provided by confidence fusion, we show the correlation between different confidence elicitation methods in Figure 2 (using WildHallu as the study case and more results are in Appendix E).
+Figure 3: Heatmaps of Spearman Correlation between different confidences in Llama3-8B-Instruct on WildHallu. Warmer colors indicate higher correlations. Atomic level: left; response level: right.
+To further explore the reasons behind the improvements provided by confidence fusion, we show the correlation between different confidence elicitation methods in Figure 3 (using WildHallu as the study case and more results are in Appendix G).
 Our findings are summarized as follows:
@@
 Confidence elicitaton methods of the same type (top left for generative and bottom right for discriminative) show stronger correlations compared to those across different types.
-This helps to explain why fusing generative and discriminative methods are effective, since these two types capture different aspects of uncertainty and are complementary to each other.
+This helps to explain why cross-category fusion strategies are effective, since these two types capture different aspects of uncertainty and are complementary to each other.
 The alignment is stronger at the response level than at the atomic level.
@@
 This highlights the need for future research on the discrepancies between generative and discriminative confidence elicitation methods, as well as how to better unify these approaches.
+Figure 4: Average confidence scores across different parts of long-form responses.
+For discriminative methods, confidence decreases as the generation progresses, while generative methods show the lowest confidence in the middle sections.
 6.2 Confidence Across Different Positions
@@
 On the contrary, in the middle parts where the models address more specific facts about individuals’ lives, careers and achievements, they tend to cover different aspects and details.
-7 Related Work
-Figure 4: An example from WildHallu dataset by Mistral-7B-Instruct.
-We only select five atomic facts for demonstration.
-The example shows the effectiveness of calculating confidence in atomic level with fusion strategy.
-Atomic Facts Generation and Verification.
-Long-form responses often contain both correct and incorrect statements, which impact the overall factuality assessments.
-Min et al.
-(2023) propose breaking long responses into atomic facts and calculating the precision of these fact pieces to determine the overall factuality score.
-Wei et al.
-(2024) and Zhao et al.
-(2024) extend this paradigm by expanding the dataset to include more domains beyond biographies.
-Song et al.
-(2024) design VeriScore for diverse long-form generation tasks that feature both verifiable and unverifiable content.
-Chiang and Lee (2024) introduce D-FActScore, specifically designed for content with ambiguous entities.
-Overall, the approach of breaking long-form responses into atomic facts for verification is nowadays widely accepted in factuality assessment.
-Uncertainty and Calibration in Long-form Generations.
-Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024).
-However, limited work has explored calibration for long-form generations.
-Huang et al.
-(2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses.
-Band et al.
-(2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation.
-Zhang et al.
-(2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs.
-None of the aforementioned studies have systematically examined calibration at a fine-grained level.
-Our work aims to fill this gap with a focus on factuality.
+6.3 The Utilities of Atomic Calibration
+While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer.
+Among these applications, we observe consistent improvements in factuality with atomic-level examination.
+Detailed experimental settings and results can be found in Appendix C.
+It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score.
+Instead, it relies entirely on the confidence of atomic claims to select and combine the most accurate claims.
+This means the final answer may contain claims from different sampled answers.
+More importantly, we observe that models with better atomic-level calibration (e.g., Qwen2 in Table 7, Appendix C) exhibit greater improvements after the reunion process, emphasizing the importance of examining and refining atomic calibration.
 7 Conclusion
-We introduce atomic calibration, a fine-grained approach for evaluating LLM calibration at the atomic claim level, addressing the limitations of traditional response-level calibration in long-form generation.
-Our experiments show that atomic calibration is well-suited for long-form generation and complements macro calibration.
-We also demonstrated that combining different confidence elicitation methods leads to better calibration results and atomic calibration can provide deeper insights into model confidence during the generation process.
-This approach offers a more reliable way to assess and improve the trustworthiness of LLM outputs in real-world applications.
+Our main contributions are three-fold: (1) We systematically study atomic calibration, which evaluates confidence calibration at the level of individual atomic claims.
+Our experiments reveal that models that appear well-calibrated at the response level perform poorly at the atomic level.
+(2) To analyze confidence elicitation methods, we categorize them into discriminative and generative methods.
+We also propose two novel fusion strategies to combine the confidence scores based on confidence agreement.
+(3) Our atomic-level analysis provides further insights into confidence methods alignment and confidence changes during generation.
+We find with discriminative methods, models show decreasing confidence in atomic facts as generation progresses.
+In contrast, generative methods show the lowest average confidence in the middle of the generation.
+Last but not least, we demonstrate the utilities of atomic calibration and propose for future research on more fine-grained confidence in long-form generation.
 Limitation
@@
 If logits are accessible, our methods can be directly applied to closed-source models without affecting the atomic calibration process.
+Third, in this work, we mainly focus on exploring different confidence elicitation methods and therefore do not apply post-hoc calibration techniques such as histogram binning or temperature scaling.
+Applying these methods makes it difficult to disentangle improvements due purely to elicitation from those due to recalibration.
+To isolate the contribution of our elicitation designs and to avoid conflating them with downstream post-processing effects, we report raw atomic and macro confidence scores, leaving a systematic study of post-hoc techniques to future work.
 Ethics Statement
@@
 where $o_{i}$ is the true fraction of positive instances in bin $i$, $e_{i}$ is the mean of the post-calibrated probabilities for the instances in bin $i$, and $P(i)$ is the empirical probability (fraction) of all instances that fall into bin $i$.
-The lower the $ECE$ value is, the better a model is calibrated.
+The lower the $ECE$ value, the better a model is calibrated.
+When labels are continuous values between 0 and 1, the ECE formulation can be generalized.
+Instead of binning instances based on binary outcomes, the continuous predictions are grouped into bins according to their predicted probability values.
+Specifically, the observed calibration error $o_{i}$ in each bin is the average of the continuous label values for the instances in that bin, and $e_{i}$ is the mean predicted probability for those instances.
+This ensures that the calibration error accounts for all possible real-valued outcomes within the range [0,1], providing a more nuanced measure of calibration when the labels are continuous.
 Brier Score
 The Brier score measures the accuracy of probabilistic predictions.
-For unidimensional predictions, it is strictly equivalent to the mean squared error as applied to predicted probabilities.
-Suppose that on each of the $n$ occasions an event can occur in only one of $r$ possible classes or categories and on one such occasion $i$, the forecast probabilities are $f_{i1},f_{i2}$, $\ldots f_{ir}$, that the event will occur in classes $1,2,\ldots r$, respectively.
-The $r$ classes are chosen to be mutually exclusive and exhaustive so that
-$\sum_{i=1}^{r}f_{ij}=1,i=1,2,3,\ldots n.$
-The Brier Score $P$ is then defined by
-$P=\frac{1}{n}\sum_{j=1}^{r}\sum_{i=1}^{n}\left(f_{ij}-E_{ij}\right)^{2},$
-where $E_{ij}$ takes the value 1 or 0 according to whether the event occurred in class $j$ or not.
+In binary classification, it compares the predicted probability of the positive class with the actual binary outcome (0 or 1).
+The Brier score is defined as:
+$P=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2},$
+where $\hat{y}_{i}$ is the predicted probability for instance $i$ and $y_{i}\in\{0,1\}$ is the actual binary outcome.
+For continuous labels in the range $[0,1]$, the Brier score can still be used, where $y_{i}$ is now a continuous value between 0 and 1.
+In this case, the Brier score becomes equivalent to the mean squared error (MSE) between predicted probabilities and the true values, and minimizing the Brier score for continuous labels is analogous to minimizing MSE.
+Both metrics aim to reduce the squared differences between predicted and true values, with lower scores indicating better calibration and accuracy.
 AUROC
@@
 Higher scores are better for AUROC, and perfect confidence score is 1, while a random confidence measure would be 0.5.
-Appendix B Prompts
+Spearman Correlation
+Following Zhang et al.
+(2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores.
+Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
+Appendix B Statistics in Atomic and Macro Calibration
+To assess the confidence of a model, we generate responses using various questions (e.g., $N$ questions).
+For each response, a single confidence score is too coarse-grained.
+Instead, we evaluate the confidence of each atomic claim, with an average of $M$ atomic claims per response.
+These individual confidences are then aggregated into a response-level confidence score.
+Atomic calibration is computed over $MN$ data points, where $M$ is the average number of atomic claims per response, and $N$ is the number of responses.
+In contrast, response-level calibration is based on $N$ data points.
+This distinction highlights the trustworthiness of the model’s confidence at both the atomic and response levels, providing a more granular view of its performance.
+From the above discussion, it follows that to ensure sufficient data points for atomic calibration, $MN$ must be large.
+In our datasets, $N$ typically exceeds 1,000, ensuring that $MN$ remains robust even when some responses have only a few calims.
+The detailed generation statistics are further illustrated in Figures 5, 6, and 7.
+Figure 5 presents the average answer length, while Figure 6 shows the average number of atomic claims per answer.
+Finally, Figure 7 highlights the percentage of answers containing fewer than 10 atomic facts.
+Figure 5: Average answer length (in words) for different models on Bios, longfact, and wildhallu.
+Figure 6: Average number of atomic claims per answer for different models on Bios, longfact, and wildhallu.
+Figure 7: Answer with less than 10 atomic facts (%) by dataset and model.
+Notably, Mistral-8$\times$7B-Instruct has 6% short answers.
+Human evaluation reveals that these responses are primarily instances where the model refuses to answer.
+Appendix C Applications
+Selective Question Answering:
+Selective question answering involves setting a confidence threshold, which can be derived from a validation set, to selectively reject questions with low confidence.
+This approach aims to improve the overall factuality of the responses by eliminating potentially unreliable answers.
+Using the Bios dataset, we evaluate the performance of three models: Llama3-8B-Instruct, Mistral-7B-Instruct, and Qwen2-7B-Instruct with Dis-Gen and Semantic Entropy (SE).
+We observe an improvement in overall factuality as we gradually rejected more questions (from 0% to 10%).
+The table below illustrates this trend, highlighting the utility of Dis-Gen in identifying accurate responses and improving selective question answering.
+Our comparison between Dis-Gen and Semantic Entropy (SE) indicates that Dis-Gen brings more significant improvements in factuality, suggesting that better calibration methods can substantially enhance the results of selective question answering.
+LLM Ensemble:
+In the LLM Ensemble method, we use three models to generate answers to the same question and select the response from the model with the highest confidence.
+This approach aims to enhance factuality by leveraging the strengths of each model.
+The Answer Distribution (AD) shows the proportion of the final response contributed by each model, highlighting the benefit of ensemble methods.
+The table below presents the results of applying this method on the Bios and WildHallu datasets, comparing two different selection strategies: DIS-Gen and SE.
+The results demonstrate that the ensemble method with DIS-Gen significantly improves factuality.
+For instance, in the Bios dataset, the factuality score increases from 0.475 to 0.556, and in the WildHallu dataset, it increases from 0.655 to 0.752.
+In contrast, using SE results in no improvements, with factuality scores even lower than the best individual model (0.484 vs 0.502 and 0.671 vs 0.701) for the Bios and WildHallu datasets, respectively.
+These findings suggest that ensembling does not always guarantee better results, and the selection strategy, such as DIS-Gen, plays a crucial role in improving factuality.
+Atomic Reunion:
+In Atomic Reunion, for each question, we begin by sampling the model’s output five times (this is also what we need to calculate Gen-Dis.
+These outputs are then broken down into atomic claims, which are individual, verifiable statements.
+Each claim is evaluated for confidence, and only those with a high confidence level are retained.
+Subsequently, we prompt a LLM, such as GPT-4o, to reassemble the selected atomic claims into a cohesive and factually accurate response.
+This method seeks to enhance factuality by utilizing smaller, more manageable pieces of information, allowing the LLM to generate a more reliable answer by combining only high-confidence atomic claims.
+The table below presents the factuality scores of the new answers generated through the Atomic Reunion approach.
+As observed, this approach leads to a significant improvement in factuality compared to the baseline models.
+Refuse Rate Llama3-8B-Instruct Mistral-7B-Instruct Qwen2-7B-Instruct
+SE
+0% 0.475 0.403 0.502
+5% 0.479 0.407 0.506
+7.5% 0.483 0.411 0.511
+10% 0.485 0.416 0.518
+DIS-Gen
+0% 0.475 0.403 0.502
+5% 0.496 0.419 0.517
+7.5% 0.511 0.438 0.533
+10% 0.528 0.465 0.557
+Table 5: Factuality Scores with Varying Refuse Rates for Selective Question Answering
+Model Bios WildHallu
+Factuality Scores Answer Distribution Factuality Scores Answer Distribution
+DIS-Gen
+Llama3-8B-Instruct 0.475 31% 0.655 29%
+Mistral-7B-Instruct 0.403 23% 0.631 27%
+Qwen2-7B-Instruct 0.502 46% 0.701 44%
+Ensemble 0.556 / 0.752 /
+SE
+Llama3-8B-Instruct 0.475 18% 0.655 15%
+Mistral-7B-Instruct 0.403 37% 0.631 44%
+Qwen2-7B-Instruct 0.502 45% 0.701 41%
+Ensemble 0.484 / 0.671 /
+Table 6: LLM Ensemble Results on Bios and WildHallu Datasets
+Model Before Atomic Reunion After Atomic Reunion
+DIS-Gen
+Llama3-8B-Instruct 0.475 0.501
+Mistral-7B-Instruct 0.403 0.441
+Qwen2-7B-Instruct 0.502 0.575
+Table 7: Factuality Scores Before and After Atomic Reunion using DIS-Gen
+Appendix D Reliability of Atomic Facts Generation and Verification
+The processes of atomic fact generation and verification have been extensively studied and validated in prior work Min et al.
+(2023); Wei et al.
+(2024); Zhao et al.
+(2024).
+For instance, FActScore Min et al.
+(2023) reports an error rate of 2%.
+In this work, we leverage their pipeline while employing stronger models from GPT-3.5 and GPT-4o, to further enhance performance.
+We, the authors, conducted additional tests comparing GPT’s atomic decompositions with ground-truth manual segmentations.
+We manually selected 30 samples for this evaluation.
+The results of our assessment are as follows:
+  • Consistency: Over 10 trials, GPT demonstrated a high inter-run consistency of 95%, indicating stable and repeatable outcomes.
+  • Error Rate: The error rate, which includes missing or overly segmented claims and misclassification of factuality, was measured at 6.4%.
+This error rate is manageable within the context of our calibration framework, suggesting that the model’s atomic fact generation is reliable.
+Appendix E Experiment Details
+We use vLLM (Kwon et al., 2023) for our LLM inference tasks, with the following parameters: temperature = 1, top-$p$ = 0.95, and a maximum output of 512 tokens.
+For discriminative confidence elicitation methods, we set the temperature to 0 and only consider the top 10 logits.
+For generative methods, we use $N=20$ samples.
+The experiments are conducted on A100-SXM-40GB GPUs.
+Running the discriminative methods takes 30 minutes for 500 samples, while the generative methods take 1.3 hours for the same number of samples.
+We use GPT-4o as the auxiliary model for generating atomic claims and fact-checking the LLM.
+Appendix F Confidence Fusion Results
+Bios LongFact WildHallu
+ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
+Gen-Binary 13.7 19.0 81.9 8.4 11.5 80.1 12.7 17.0 81.3
+Dis-Rating 44.5 42.5 65.0 10.0 14.2 67.9 19.7 23.9 68.1
+Dis-Context 24.8 26.0 77.5 15.7 16.1 75.3 20.6 21.7 79.8
+[…]
+HMean 14.3 18.3 82.1 7.6 12.2 81.0 11.5 16.6 83.4
+ProdConf 14.2 18.3 82.3 9.5 13.0 81.0 7.9 15.9 83.5
+WAvg 10.6 16.6 84.7 5.5 10.7 82.1 12.3 16.4 84.4
+[…]
+DampedFusion 10.2 16.5 84.6 5.9 10.9 81.9 7.1 15.4 83.8
+Table 8: Atomic calibration results of confidence fusion strategies for Mistral-7B-Instruct. The fusion results are based on Gen-Binary and Dis-Context.
+Bios LongFact WildHallu
+ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
+Gen-Binary 10.9 16.7 83.8 6.3 9.9 81.9 9.5 14.0 82.5
+Dis-Rating 41.5 39.7 64.2 3.5 11.7 62.6 8.2 18.1 70.4
+Dis-Context 26.5 28.3 75.5 13.9 14.8 77.9 17.2 19.4 81.2
+[…]
+HMean 11.1 16.9 82.7 2.7 9.6 81.7 4.9 13.6 83.8
+ProdConf 12.4 17.0 83.1 8.3 10.6 81.7 7.2 13.7 83.9
+WAvg 10.7 15.9 84.8 2.6 9.2 82.8 6.8 13.5 84.3
+[…]
+DampedFusion 10.2 15.8 84.5 2.6 9.3 82.9 5.2 13.3 84.4
+Table 9: Atomic calibration results of confidence fusion strategies for Qwen2-7B-Instruct. The fusion results are based on Gen-Binary and Dis-Context.
+Appendix G Confidence Alignment
+(a) Mistral-7B-Instruct : atomic
+(b) Mistral-7B-Instruct : response
+(c) Qwen2-7B-Instruct : atomic
+(d) Qwen2-7B-Instruct : response
+(e) Llama3-8B-Instruct : atomic
+(f) Llama3-8B-Instruct : response
+Figure 8: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for Bios.
+Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
+(a) Mistral-7B-Instruct : atomic
+(b) Mistral-7B-Instruct : response
+(c) Qwen2-7B-Instruct : atomic
+(d) Qwen2-7B-Instruct : response
+(e) Llama3-8B-Instruct : atomic
+(f) Llama3-8B-Instruct : response
+Figure 9: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for LongFact.
+Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
+(a) Mistral-7B-Instruct : atomic
+(b) Mistral-7B-Instruct : response
+(c) Qwen2-7B-Instruct : atomic
+(d) Qwen2-7B-Instruct : response
+(e) Llama3-8B-Instruct : atomic
+(f) Llama3-8B-Instruct : response
+Figure 10: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for WildHallu.
+Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
+Appendix H Case Study
+Figure 11: An example from WildHallu dataset by Mistral-7B-Instruct.
+We only select five atomic facts for demonstration.
+The example shows the effectiveness of calculating confidence in atomic level with fusion strategy.
+Appendix I Prompts
 Prompts to Generate Responses Bios: Tell me a bio of <ENTITY>.
@@
 Table 15: Prompt for Dis-Rating.
-Appendix C Experiment Details
-We use vLLM (Kwon et al., 2023) for our LLM inference tasks, with the following parameters: temperature = 1, top-$p$ = 0.95, and a maximum output of 512 tokens.
-For discriminative confidence elicitation methods, we set the temperature to 0 and only consider the top 10 logits.
-For generative methods, we use $N=20$ samples.
-The experiments are conducted on A100-SXM-40GB GPUs.
-Running the discriminative methods takes 30 minutes for 500 samples, while the generative methods take 1.3 hours for the same number of samples.
-We use GPT-4o as the auxiliary model for generating atomic claims and fact-checking the LLM.
-Appendix D Confidence Fusion Results
-Bios LongFact WildHallu
-SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
-Gen-Binary 79.8 5.6 5.6 52.7 3.0 3.0 70.0 7.2 7.8
-Dis-Rating 73.8 25.7 25.7 34.1 2.8 2.9 71.7 3.8 3.6
-Dis-Context 55.4 33.2 34.0 29.7 5.4 5.6 65.9 9.1 9.5
-[…]
-HMean 81.7 11.5 11.5 54.5 2.0 2.0 78.7 3.2 3.0
-ProdConf 82.2 7.7 7.1 54.9 14.2 14.3 78.7 12.2 12.1
-WAvg 81.8 10.9 11.0 55.6 1.5 0.9 77.2 5.0 5.0
-Table 11: Macro calibration results of confidence fusion strategies for Llama3-8B-Instruct.
-Bios LongFact WildHallu
-SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
-Gen-Binary 74.9 8.0 8.5 64.1 2.5 2.5 73.9 10.3 10.3
-Dis-Rating 55.0 41.5 41.4 40.8 4.2 4.4 60.4 16.7 16.9
-Dis-Context 79.7 8.2 8.3 47.9 4.2 4.1 72.3 5.7 6.1
-[…]
-HMean 83.5 5.8 6.2 62.5 1.9 2.1 77.9 6.5 6.6
-ProdConf 83.8 13.1 13.0 62.5 10.6 10.4 78.0 7.2 7.1
-WAvg 82.0 8.0 8.0 65.1 1.8 2.0 78.4 8.9 8.9
-Table 12: Macro calibration results of confidence fusion strategies for Mistral-7B-Instruct.
-Bios LongFact WildHallu
-SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$ SC $\uparrow$ UCCE $\downarrow$ QCCE $\downarrow$
-Gen-Binary 72.4 5.7 5.4 67.6 1.8 2.0 72.2 6.4 6.5
-Dis-Rating 63.0 41.0 40.7 29.9 4.4 4.4 54.0 8.9 9.1
-Dis-Context 66.5 14.4 14.8 40.6 3.9 3.9 66.8 3.7 4.0
-[…]
-HMean 76.2 7.8 8.0 62.4 1.4 1.4 74.7 2.3 2.2
-ProdConf 76.0 12.9 12.6 62.5 10.4 10.4 74.9 12.7 12.7
-WAvg 75.7 7.5 8.4 66.4 1.2 1.4 75.6 4.5 4.3
-Table 13: Macro calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
-Bios LongFact WildHallu
-ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
-Gen-Binary 13.7 19.0 81.9 8.4 11.5 80.1 12.7 17.0 81.3
-Dis-Rating 44.5 42.5 65.0 10.0 14.2 67.9 19.7 23.9 68.1
-Dis-Context 24.8 26.0 77.5 15.7 16.1 75.3 20.6 21.7 79.8
-[…]
-HMean 14.3 18.3 82.1 7.6 12.2 81.0 11.5 16.6 83.4
-ProdConf 14.2 18.3 82.3 9.5 13.0 81.0 7.9 15.9 83.5
-WAvg 10.6 16.6 84.7 5.5 10.7 82.1 12.3 16.4 84.4
-Table 14: Atomic calibration results of confidence fusion strategies for Mistral-7B-Instruct.
-Bios LongFact WildHallu
-ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$ ECE $\downarrow$ BS $\downarrow$ AUROC $\uparrow$
-Gen-Binary 10.9 16.7 83.8 6.3 9.9 81.9 9.5 14.0 82.5
-Dis-Rating 41.5 39.7 64.2 3.5 11.7 62.6 8.2 18.1 70.4
-Dis-Context 26.5 28.3 75.5 13.9 14.8 77.9 17.2 19.4 81.2
-[…]
-HMean 11.1 16.9 82.7 2.7 9.6 81.7 4.9 13.6 83.8
-ProdConf 12.4 17.0 83.1 8.3 10.6 81.7 7.2 13.7 83.9
-WAvg 10.7 15.9 84.8 2.6 9.2 82.8 6.8 13.5 84.3
-Table 15: Atomic calibration results of confidence fusion strategies for Qwen2-7B-Instruct.
-Appendix E Confidence Alignment
-(a) Mistral-7B-Instruct : atomic
-(b) Mistral-7B-Instruct : response
-(c) Qwen2-7B-Instruct : atomic
-(d) Qwen2-7B-Instruct : response
-(e) Llama3-8B-Instruct : atomic
-(f) Llama3-8B-Instruct : response
-Figure 5: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for Bios.
-Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
-(a) Mistral-7B-Instruct : atomic
-(b) Mistral-7B-Instruct : response
-(c) Qwen2-7B-Instruct : atomic
-(d) Qwen2-7B-Instruct : response
-(e) Llama3-8B-Instruct : atomic
-(f) Llama3-8B-Instruct : response
-Figure 6: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for LongFact.
-Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.
-(a) Mistral-7B-Instruct : atomic
-(b) Mistral-7B-Instruct : response
-(c) Qwen2-7B-Instruct : atomic
-(d) Qwen2-7B-Instruct : response
-(e) Llama3-8B-Instruct : atomic
-(f) Llama3-8B-Instruct : response
-Figure 7: Heatmaps comparing the Spearman correlation between generative and discriminative confidence elicitation methods for WildHallu.
-Results shown for Mistral-7B-Instruct, Qwen2-7B-Instruct, and Llama3-8B-Instruct.

### v3

@@
 While large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023; OpenAI, 2022) excel in various tasks, they still struggle with trustworthiness issues.
-LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023).
-To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017).
+LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2024a; Shelmanov et al., 2025; He et al., 2025), which limits their application in high-risk real-world scenarios (Hu et al., 2023).
+To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017; Zhang et al., 2025a, b).
+In this work, we distinguish between calibration as a property and as a process.
+We use the term atomic calibration to refer to the property of each atomic claim’s confidence being well-aligned with its factuality (Zhou et al., 2025).
+In contrast, the extra post-hoc methods used to achieve this alignment (such as temperature scaling and platt scaling) are referred to as calibration methods or a calibrator, which is not considered in this paper.
 A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024).
@@
 Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024), using datasets like TriviaQA and Natural Questions (Joshi et al., 2017), where answers are typically under 10 words.
-In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024; Yang et al., 2024b), spanning hundreds or thousands of words.
+In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024b; Yang et al., 2025a, b), spanning hundreds or thousands of words.
 In such cases, response quality is not simply binary, as answers may mix accurate and inaccurate statements.
-Recent work has begun to address calibration in long-form generation (Zhang et al., 2024; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024).
+Recent work has begun to address calibration in long-form generation (Zhang et al., 2024b; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024).
 Several approaches estimate a single confidence score for the entire response (macro calibration; upper, Figure 1), while others assess confidence at the level of atomic claims (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024) (atomic calibration; lower, Figure 1).
@@
 Zhang et al.
-(2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs.
+(2024b) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs.
 Another line of work (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024; Yuan et al., 2024) decomposes sentences into atomic claims and assigns confidence scores to each claim.
@@
 $\mathbb{P}(y\mid f(x)_{y}=\beta)=\beta$
-⇏$\not\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta$
-⇏$\not\Rightarrow$ $\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta,\forall i\in\{1,....,N\}.$
+$\not\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta$
+$\not\Rightarrow$ $\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta,\forall i\in\{1,..,N\}.$
 4 Confidence Elicitation Methods
@@
 4.1 Generative Methods
-Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024; Jiang et al., 2024).
+Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024b; Jiang et al., 2024).
 Generally, an additional natural language inference (NLI) model is used to calculate the consistency.
@@
 For example, when $\alpha=0.5$, confidences of 0 and 1 are treated the same as confidences of 0.4 and 0.6, although the former may indicate higher uncertainty due to a larger disagreement.
-To address this, we propose two simple but effective methods that consider confidence disagreement $d=B-A$.
+To address this, we propose two simple but effective methods that consider confidence disagreement d=BAd=B-A.
 AdjustedAlpha adjusts the weight $\alpha$ based on the confidence difference:
@@
 Models.
-We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024a).
+We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024).
 Datasets.
@@
 EigV 36.8 21.2 43.0 13.0 3.2 23.8 17.2 11.3 53.0
-[…]
+Dis-Context 34.0 17.3 55.4 5.6 1.9 29.7 9.5 4.8 65.9
 Dis-Rating 25.7 11.7 73.8 2.9 1.6 34.1 3.6 3.5 71.7
@@
 EigV 35.9 23.3 45.3 13.3 3.5 36.5 16.9 13.9 54.4
-[…]
+Dis-Context 8.3 3.4 79.7 4.1 1.4 47.9 6.1 4.3 72.3
 Dis-Rating 41.4 22.7 55.0 4.4 1.7 40.8 16.9 9.8 60.4
@@
 EigV 22.4 22.5 50.7 6.2 3.5 39.8 8.5 13.5 50.2
-[…]
+Dis-Context 14.8 5.8 66.5 3.9 1.7 40.6 4.0 3.5 66.8
 Dis-Rating 40.7 21.3 63.0 4.4 1.9 29.9 9.1 6.3 54.0
@@
 Llama3-70B-Instruct 10.0 16.5 82.5 8.3 9.3 73.7 9.5 12.3 78.3
-[…]
+Mistral-7B-Instruct 13.7 19.0 81.9 8.4 11.5 80.1 12.7 17.0 81.3
 Mistral-8x7B-Instruct 12.3 18.5 79.8 7.8 9.0 76.3 9.8 13.4 77.8
-[…]
+Qwen2-7B-Instruct 10.9 16.7 83.8 6.3 9.9 81.9 9.5 14.0 82.5
 Qwen2-57B-Instruct 10.5 18.1 82.3 7.8 10.0 78.3 9.2 13.6 81.7
@@
 Llama3-70B-Instruct 10.6 19.3 73.2 4.2 8.0 74.2 4.3 11.5 81.2
-[…]
+Mistral-7B-Instruct 44.5 42.5 65.0 10.0 14.2 67.9 19.7 23.9 68.1
 Mistral-8x7B-Instruct 15.3 22.6 70.8 5.3 8.6 72.6 7.6 14.7 72.9
-[…]
+Qwen2-7B-Instruct 41.5 39.7 64.2 3.5 11.7 62.6 8.2 18.1 70.4
 Qwen2-57B-Instruct 23.2 27.0 69.3 2.2 9.8 71.3 5.2 15.2 77.2
@@
 Dis-Context 35.5 35.8 74.5 11.9 13.6 74.4 12.5 16.5 83.5
-[…]
+MinConf 6.2 17.1 83.2 10.7 12.2 77.4 9.0 13.9 85.8
 HMean 9.8 17.4 84.0 4.1 11.0 79.6 5.6 13.3 87.0
@@
 WAvg 10.9 17.4 84.4 3.3 10.3 79.9 5.1 13.0 87.0
-[…]
+AdjustedAlpha 4.1 15.8 85.2 3.4 10.2 80.4 4.3 12.6 88.3
 DampedFusion 5.0 15.6 84.7 3.5 9.8 80.0 4.8 12.4 87.9
@@
 6.3 The Utilities of Atomic Calibration
-While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer.
+While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024b), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer.
 Among these applications, we observe consistent improvements in factuality with atomic-level examination.
 Detailed experimental settings and results can be found in Appendix C.
-It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score.
+It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024b) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score.
 Instead, it relies entirely on the confidence of atomic claims to select and combine the most accurate claims.
@@
 Following Zhang et al.
-(2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores.
+(2024b), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores.
 Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
@@
 Figure 6: Average number of atomic claims per answer for different models on Bios, longfact, and wildhallu.
-Figure 7: Answer with less than 10 atomic facts (%) by dataset and model.
-Notably, Mistral-8$\times$7B-Instruct has 6% short answers.
-Human evaluation reveals that these responses are primarily instances where the model refuses to answer.
+[…]
 Appendix C Applications
@@
 Dis-Context 24.8 26.0 77.5 15.7 16.1 75.3 20.6 21.7 79.8
-[…]
+MinConf 14.1 18.3 82.0 8.6 12.7 80.7 7.6 16.0 83.2
 HMean 14.3 18.3 82.1 7.6 12.2 81.0 11.5 16.6 83.4
@@
 WAvg 10.6 16.6 84.7 5.5 10.7 82.1 12.3 16.4 84.4
-[…]
+AdjustedAlpha 9.8 16.7 85.0 5.8 10.5 81.8 6.5 15.2 84.0
 DampedFusion 10.2 16.5 84.6 5.9 10.9 81.9 7.1 15.4 83.8
@@
 Dis-Context 26.5 28.3 75.5 13.9 14.8 77.9 17.2 19.4 81.2
-[…]
+MinConf 11.3 16.9 82.4 6.3 10.3 80.5 5.0 13.8 82.6
 HMean 11.1 16.9 82.7 2.7 9.6 81.7 4.9 13.6 83.8
@@
 WAvg 10.7 15.9 84.8 2.6 9.2 82.8 6.8 13.5 84.3
-[…]
+AdjustedAlpha 8.9 16.0 84.6 2.9 9.1 82.6 4.5 13.2 84.1
 DampedFusion 10.2 15.8 84.5 2.6 9.3 82.9 5.2 13.3 84.4